OpenStudy (anonymous):

Simplify: 3^4 + 2 * 3^4

OpenStudy (anonymous):

solve first 3^4, then multiply by 2, then add. 3^4+2*3^4=81+2*81=81+162=243

OpenStudy (anonymous):

the answer is 243

geerky42 (geerky42):

Factor \(3^4\) out and you should have \(3^4(1+2) = 3^4\cdot3 = \boxed{3^5}\)

geerky42 (geerky42):

Again, \(3^5 = 243\)